An ideal Stirling cycle cooler, using helium (considered ideal) as its working fluid, has a compression temperature Tc=32°C and an expansion temperature
has a compression temperature Tc=32°C and an expansion temperature
TE=-21°C.
If state 1, at the beginning of the compression, is defined with a volume V1=40
cm3 and a pressure P1=1.8 MPa, and state 2 is defined with a volume V2=33
cm3, what is the heat qE absorbed during the expansion process (in Joules)?
1
Expert's answer
2020-08-14T07:30:37-0400

Given data are

TH = 320 C + 273 = 305 k

TL = -210 C + 273 = 252 k

V1 = 40 cm3 , V3 = 33 cm3

n = ?

work done in isothermal expansion

w = nRT ln v2/v1

= 2.303 *1 *8.31 * 305 log (40/33)

= 487.313 Joules

in isothermal process change in u = 0

from its law of thermodynamics

dq = change in u + change in w

dq = 0 + w

hence,

heat absorbed q = w

q = 487.313 Joules